On January 1, 2017, Citywide Sales issued $23,000 in bonds for $30,800. These are eight-year bonds with a stated rate of 13% and pay semiannual interest. Citywide Sales uses the straight-line method to amortize the bond premium. On June 30, 2017, when Citywide makes the first payment to bondholders, what is the amount that will be reported as Interest Expense? (Round your intermediate answers to the nearest dollar.)
Answer #1

Premium on bonds payable = 30800-23000 = 7800

Semiannually premium amortization = 7800/16 = 487.50 or 488

Interest expense = Interest paid - Premium amortization

= (23000*13%*6/12)-488

Interest expense = 1007

So 1007 will be reported as interest expense.